OpenStudy (anonymous):

Show that cos2x-cos4x/sin2x+sin4x= tanx

OpenStudy (***[isuru]***):

hi, r u familiar with these ? \[\cos A - cosB = 2\sin \frac{ A+B }{2 }\sin \frac{ B-A }{2 }\] \[sinA + sinB = 2\sin\frac{A+B }{ 2 }\cos \frac{ A-B }{2 }\]

OpenStudy (anonymous):

Oh yes! Thank u so much!:)

OpenStudy (***[isuru]***):

u r welcome!!